OpenStudy (anonymous):

Find the first 5 terms of the sequence: aa = 2n - 5. Enter your answers separated by commas; example: -1, 2, -3, 4, -5.

OpenStudy (haseeb96):

n is the natural number so just put n=1,2,3,,4,5 in this equation a = 2n-5 for n=1 a1=2-5 =-3 a2=2(2)-5=4-5=-1 a3=2(3)-5= 6-5=1 a4=2(4)-5= 8-5=3 a5=2(5)-5= 10-5=5 the first five terms are given below -3,-1,1,3,5 this is the correct answer

OpenStudy (haseeb96):

i hope u got it now ...

OpenStudy (haseeb96):

i know i give u direct answer which is against the Code of conduct of OS but idk how to deal with this problem

OpenStudy (anonymous):

hmm it says its wrong

OpenStudy (anonymous):

@Haseeb96

OpenStudy (haseeb96):

My answers are correct absolutely

OpenStudy (anonymous):

well its weird my online class says its rong

OpenStudy (haseeb96):

no it is correct i m sure

OpenStudy (anonymous):

hmm its -3,-1,1,3,5 but thats for the equation an and this one is aa

OpenStudy (anonymous):

Find the first 5 terms of the sequence: aa = 2n - 5.

OpenStudy (haseeb96):

what is aa >?

geerky42 (geerky42):

\(a_a=2n-5\) Or do you mean \(a_n = 2n-5\)?

geerky42 (geerky42):

? @fani1996

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

aa=2n-5

geerky42 (geerky42):

Well, in that case, I guess sequence goes like this: \(2n-5,~2n-5,~2n-5,~2n-5,~2n-5\) Can you tell us what was recently discussed in your class or something, so we can have idea what we are working with?

OpenStudy (anonymous):

well this is series and summatation notation

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@myininaya

OpenStudy (haseeb96):

@uri please have little eye on her question

OpenStudy (camper4834):

if you are starting with 1 then the first term is -3 then plug -3 in to get the next. -3, -11, -27, -59, -123

OpenStudy (camper4834):

that is if you are starting with 1...

OpenStudy (anonymous):

says its wrong

OpenStudy (anonymous):

@camper4834

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

put n=1 and tell me what u got ??

OpenStudy (anonymous):

PUT IT WHEERE

OpenStudy (gorv):

a=2n-5

OpenStudy (anonymous):

HMMM CONFUSED

OpenStudy (gorv):

we need to put value of n

OpenStudy (gorv):

for first term n=1

OpenStudy (gorv):

for second n=2

OpenStudy (gorv):

plug in n=1 and tell me what u got ??

OpenStudy (gorv):

ok

OpenStudy (gorv):

a=2*1-5=2-5=-3 it is first term

OpenStudy (gorv):

now plug in n=2 and tell what u got ??

OpenStudy (anonymous):

-1

OpenStudy (gorv):

yeah that will be second term

OpenStudy (gorv):

for third term plug in n=3

OpenStudy (gorv):

for 4th term plug in n=4 and so on

OpenStudy (gorv):

@fani1996 got me ??

OpenStudy (anonymous):

if you thi nk its -3,-1,1,3,5 ur wrong

OpenStudy (anonymous):

@HARSH123

OpenStudy (anonymous):

So if you want the first 5 terms, you want A1, A2, A3, A4, and A5. An = 2n-5 So A1 = 2(1) - 5 = -3 A2 = 2(2) - 5 = -1 A3 = 2(3) - 5 = 1 A4 = 2(4) - 5 = 3 A5 = 2(5) - 5 = 5 -3, -1, 1, 3, 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it says its wron g

OpenStudy (anonymous):

who

OpenStudy (anonymous):

so u have answers

OpenStudy (anonymous):

so u have answers